OpenStudy (anonymous):

Find the derivative of: x^3+5/x

OpenStudy (aravindg):

Derivative of x^n=n*x^{n-1}

OpenStudy (anonymous):

I tried working on this problem using the power rule and got \[3x^{2}-1x^{-2}\] as my answer, am I on the right track or not?

OpenStudy (aravindg):

First term is right. For second term you forgot to multiply by 5.

OpenStudy (anonymous):

Why? Can you explain?

OpenStudy (aravindg):

5/x=5 *(1/x) Derivative of 1/x is -1/x^2 Then multiply by 5 5*(-1/x^2}

OpenStudy (aravindg):

Note that derivative of k(f(x))=k(f'(x))

OpenStudy (anonymous):

Ok, so then my answer would be: \[3x ^{2}-5x^{-2}\] ??

OpenStudy (anonymous):

Yes

OpenStudy (aravindg):

Yes :)

OpenStudy (anonymous):

Ok, thank you very much @AravindG & @Andras I just wasn't too sure I was doing things right, glad you helped me identify my mistake

OpenStudy (aravindg):

yw :)